### v1

@@ 2.2.2 Dory-Guest-Harris instability
We also demonstrate the computation of (2.6) for the ring-shaped equilibrium as discussed in [8],
@@
A special case for this equilibrium is when $j=0$, when it recovers the Gaussian equilibrium. It was further identified [8] the equilibrium (2.28) continue being unconditionally stable for $j\in\{0,1,2\}$. However, for $j\geq 3$, the equilibrium (2.28) can be unstable, and this instability is termed the Dory-Guest-Harris instability. Intuitively, for larger $j$, the distribution becomes more concentrated on a ring [22, 23], and this concentration leads to instability. The cases of $j=3,4,5,6$ were thoroughly studied in [10].
@@
Again without loss of generality, we fix $|\mathbf{k}|^{2}=1$, i.e. $\mathbf{k}=(1,0)^{T}$ or $\mathbf{k}=(0,1)^{T}$, and numerically compute $P$ as a function of $\lambda$ (2.9). In Figure 2 we plot both the distribution function (2.29) and $P(\mathbf{k}=(1,0),\lambda)$. Clearly, in this case, we find a root with a positive real component, indicating the instability of this equilibrium.
@@ 3 Control with an external electric field
$=\hat{\chi}(\mathbf{k})\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)\,$,
@@
$\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)=\mathrm{exp}{\left(-\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})-\frac{\theta}{2}k_{3}^{2}t^{2}\right)}\,.$
@@
    $\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)\leq e^{-\frac{\theta}{2}k_{3}^{2}t^{2}}\,$,
@@
    $\mathcal{E}_{\mathbf{k}}(t)=\frac{1}{2}|\mathbf{E}_{k}|^{2}=\frac{1}{2|\mathbf{k}|^{2}}\hat{\rho}^{2}(t,\mathbf{k})\leq\frac{1}{2}|\hat{\chi}(\mathbf{k})|^{2}|\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)|^{2}\leq Ce^{-At^{2}}\,$,
@@
  • For Bernstein modes, i.e. $k_{3}=0$
    $\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)=\mathrm{exp}{\left(-\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})\right)}\leq 1\,$,
    then
    $\mathcal{E}_{\mathbf{k}}(t)\leq\frac{1}{2}|\hat{\chi}(\mathbf{k})|^{2}|\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)|^{2}\leq\frac{1}{2}|\hat{\chi}(\mathbf{k})|^{2}\,.$
    It is bounded and is periodic in time.
@@
$|\mathbf{k}|^{2}L[\hat{\Phi}]+L[\hat{S}$] $=c\mathbf{k}\cdot(I+L[\hat{\mu}\tilde{\mathbf{R}}])\mathbf{k}\,$, (3.10)
@@
$\hat{\Phi}(t,\mathbf{k})$ $=-\frac{1}{|\mathbf{k}|^{2}}\hat{f}_{0}(\mathbf{k},(\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3})+c\delta(t)+c\frac{1}{|\mathbf{k}|^{2}}\hat{\mu}(\tilde{\mathbf{R}}(t)\mathbf{k})\mathbf{k}\cdot\tilde{\mathbf{R}}(t)\mathbf{k}\,$,
@@
$\hat{U}=\frac{\hat{S}}{|\mathbf{k}|^{2}}-\frac{1}{|\mathbf{k}|^{2}}c\mathbf{k}\cdot\hat{\mu}\tilde{\mathbf{R}}\mathbf{k}\,.$
@@
$\mathbf{k}\cdot\tilde{\mathbf{R}}\mathbf{k}=\frac{\sin(B_{0}t)}{B_{0}}|\mathbf{k}_{\perp}|^{2}+k_{3}^{2}t\,.$
This structure prevents us from directly taking the Fourier inverse transform for an explicit solution for $U$, as was done in [21] for the Vlasov-Poisson system without external $\mathbf{B}$.
@@ 4 Numerical experiments
The nonlinear system is computed using a semi-Lagrangian scheme with Strang splitting [24, 25] for the time integration. This means from one step to the next, we first solve
@@
The code is written in JAX, and the computation was performed on an NVIDIA Quadro RTX 6000 GPU with 24 GB of memory.
@@ 4.1.1 Initial perturbation as a sine wave in space
When $c=0$, the corresponding electric energy and Fourier mode coefficient are presented in Figure 12(A)-(B). The electric energy is suppressed under $10^{-5}$ for all time, and the density profile has much less variation comparing Figure 11 to Figure 10.
We also tested setting $c$ to be at different values. The results are shown in Figure 12, with $c=0.001,c=0.005,c=0.01$. For all these tests, the electric energy can be controlled below the level of $10^{-5}$ for all time, thus suppressing the instability as in Figure 9. However, we also notice that in Figure 9 (G)-(H), for $c$ being $0.01$, the electric energy gradually grows after $t=300$. This indicates that the perturbation induced by the term $c\frac{\sin(B_{0}t)}{B_{0}}\hat{\mu}$ in (3.10) might lead to another instability in the system.
@@ 4.2 Gaussian equilibrium
Gaussian equilibrium is expected to be a linearly stable equilibrium state according to the linear analysis when the initial perturbation is small. We now study the situation when a large perturbation is posed on the Gaussian equilibrium. This is beyond the linear stable regime, and potential instability may be observed. This is indeed the case. We set $B_{0}=1$, adopt the Kelvin-Helmholtz type instability example in [26] and use the following initial condition:
@@ 5 Conclusion
In this paper, we study the control problem of plasma living in an external magnetic field, modeled by the magnetized Vlasov–Poisson equation. We derived the dispersion relation in this setting and formulated its associated Penrose condition (Proposition 2.2). As a case study, we investigate two equilibrium distributions: a Gaussian equilibrium and a ring-shaped equilibrium. The latter generates a positive root for the associated PDE operator, and this leads to linear instability (Dory–Guest–Harris instability), and electric energy accordingly exponentially grows in the linear regime.
Upon obtaining this analytical result, we investigate the control problem and seek to suppress the growth of electric energy by imposing an external electric field. The formula for the dispersion relation is instrumental for this purpose. We proposed two different strategies, both of which achieve suppression by eliminating positive roots. Through analysis, it is also seen that one of these strategies recovers the free-streaming solution. In this case, one can further derive an explicit formula for the electric energy and show that the electric energy stays bounded. We numerically demonstrated the effectiveness of these strategies and observed that such a strategy is valid even in a certain nonlinear regime.
[…]
@@ Acknowledgements
P.C. thanks Genia Vogman, Xiaochun Tian, and Qihao Ye for useful discussions on Dory-Guest-Harris instability and related details. R.J. is supported by the National Science Foundation under Grant No. 2409066. P.C., Q.L., and Y.Y. are supported by the National Science Foundation under Grant No.2308440.
@@
The authors have no competing interests to declare that are relevant to the content of this article. The data and codes used in this study are openly available in the GitHub repository [^2] https://github.com/Peiyi-wisc/Magnetized_VP_control.git, and archived in Zenodo [^3] https://doi.org/10.5281/zenodo.15109942.
@@ Appendix A Computation of $P(\mathbf{k},\lambda)$
here $M(\cdot,\cdot,\cdot)$ is the Kummer’s function (or the hypergeometric function), the last line is obtained in [27] together with Kummer’s transform. One specific property of Kummer’s function is that for $j$ being nonnegative integers, $M(-j,\cdot,\cdot)$ can be expressed as a polynomial of degree no more than $j$.

### v2

@@ 2.2.2 Dory-Guest-Harris instability
We also demonstrate the computation of (2.6) for the ring-shaped equilibrium as discussed in [DGH-1965-PRL],
@@
A special case for this equilibrium is when $j=0$, when it recovers the Gaussian equilibrium. It was further identified [DGH-1965-PRL] the equilibrium (2.28) continue being unconditionally stable for $j\in\{0,1,2\}$. However, for $j\geq 3$, the equilibrium (2.28) can be unstable, and this instability is termed the Dory-Guest-Harris instability. Intuitively, for larger $j$, the distribution becomes more concentrated on a ring [Tataronis-Crawford-1965-abs-instable, Tataronis_Crawford_1970], and this concentration leads to instability. The cases of $j=3,4,5,6$ were thoroughly studied in [VOGMAN2014101].
@@
Again without loss of generality, we fix $|\mathbf{k}|^{2}=1$, i.e. $\mathbf{k}=(1,0)^{\top}$ or $\mathbf{k}=(0,1)^{\top}$, and numerically compute $P$ as a function of $\lambda$ (2.9). In Figure 2 we plot both the distribution function (2.29) and $P(\mathbf{k}=(1,0),\lambda)$. Clearly, in this case, we find a root with a positive real component, indicating the instability of this equilibrium.
@@ 3 Control with an external electric field
$=\hat{\chi}(\mathbf{k})\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)\,$,
@@
$\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)=\mathrm{exp}{\left(-\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})-\frac{\theta}{2}k_{3}^{2}t^{2}\right)}\,.$
@@
    $\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)\leq e^{-\frac{\theta}{2}k_{3}^{2}t^{2}}\,$,
@@
    $\mathcal{E}_{\mathbf{k}}(t)=\frac{1}{2}|\mathbf{E}_{\mathbf{k}}|^{2}=\frac{1}{2|\mathbf{k}|^{2}}\hat{\rho}^{2}(t,\mathbf{k})\leq\frac{1}{2}|\hat{\chi}(\mathbf{k})|^{2}|\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)|^{2}\leq Ce^{-At^{2}}\,$,
@@
  • For Bernstein modes, i.e. $k_{3}=0$,
    $\hat{\mu}((\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)=\mathrm{exp}{\left(-\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})\right)}\leq 1\,$,
    then the electric energy is periodic and bounded:
    $\mathcal{E}_{\mathbf{k}}(t)$ $=\frac{1}{2|\mathbf{k}_{\perp}|^{2}}|\hat{\chi}(\mathbf{k})|^{2}\mathrm{exp}{\left(-2\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})\right)}\leq\frac{1}{2}|\hat{\chi}(\mathbf{k})|^{2}\,.$
    The periodicity is shown in the $\cos$ function in the exponential factor, with the period being $\frac{2\pi}{B_{0}}$.
@@
To provide an additional concrete realization of the general control design, we propose this section control strategy to set $c$ to be a constant in (3.6). Allowing $c$ to be an arbitrary constant gives a greater flexibility for exercising control.
@@
$|\mathbf{k}|^{2}L[\hat{\Phi}]+L[\hat{S}$] $=c\mathbf{k}\cdot(I+L[\hat{\mu}\tilde{\mathbf{R}}])\mathbf{k}\,.$ (3.10)
@@
$\hat{\Phi}(t,\mathbf{k})$ $=-\frac{1}{|\mathbf{k}|^{2}}\hat{f}_{0}(\mathbf{k},(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3})+c\delta(t)+c\frac{1}{|\mathbf{k}|^{2}}\hat{\mu}(\tilde{\mathbf{R}}(t)\mathbf{k})\mathbf{k}\cdot\tilde{\mathbf{R}}(t)\mathbf{k}\,$,
@@
$\hat{U}=\frac{\hat{S}}{|\mathbf{k}|^{2}}-\frac{1}{|\mathbf{k}|^{2}}c\mathbf{k}\cdot\hat{\mu}\tilde{\mathbf{R}}\mathbf{k}\,.$ (3.11)
@@
$\mathbf{k}\cdot\tilde{\mathbf{R}}\mathbf{k}=\frac{\sin(B_{0}t)}{B_{0}}|\mathbf{k}_{\perp}|^{2}+k_{3}^{2}t\,.$ (3.12)
This structure prevents us from directly taking the Fourier inverse transform for an explicit solution for $U$, as was done in [EINKEMMER-Li-Mouhot-Yue-2025] for the Vlasov-Poisson system without external $\mathbf{B}$.
The performance of this control is not as clean as that of Control Strategy 1. Since it is not able to recover the free-streaming solution, this controlled system maintains nonlinear features, and the performance needs to be discussed case-by-case. Recall the definition of the electric energy of the $\mathbf{k}$-th mode ($\mathcal{E}_{\mathbf{k}}(t)=\frac{1}{2}|\mathbf{k}|^{2}|\hat{U}_{\mathbf{k}}|^{2}$) and the Poisson equation (2.4), following (3.11), we have:
$\mathcal{E}_{\mathbf{k}}(t)$ $=\frac{1}{2}|\mathbf{k}|^{2}|\hat{U}_{\mathbf{k}}|^{2}=\frac{1}{2}\left|\hat{S}-c\mathbf{k}\cdot\hat{\mu}\tilde{\mathbf{R}}\mathbf{k}\right|^{2}\leq|\hat{S}|^{2}+c^{2}\left|\mathbf{k}\cdot\hat{\mu}\tilde{\mathbf{R}}\mathbf{k}\right|^{2}$
$\leq|\hat{f}_{0}(\mathbf{k},(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3})|^{2}+c^{2}|\hat{\mu}|^{2}\left|\frac{1}{B_{0}}|\mathbf{k}_{\perp}|^{2}+k_{3}^{2}t\right|^{2}\,.$
where we used the definition of $S$, (2.16) and (3.12). If the initial profile is assumed to be of a factored form as in Theorem 3.1 with $\mu$ being a Gaussian, we obtain the following estimate of the temporal behavior of the electric energy:
$\mathcal{E}_{\mathbf{k}}(t)$ $\leq\mathrm{exp}{\left(-2\theta(1-\cos(B_{0}t))(k_{1}^{2}+k_{2}^{2})-\frac{\theta}{2}k_{3}^{2}t^{2}\right)}\left(|\chi(\mathbf{k})|^{2}+c^{2}\left|\frac{1}{B_{0}}|\mathbf{k}_{\perp}|^{2}+k_{3}^{2}t\right|^{2}\right)\,.$
That is, for non-Bernstein modes, $k_{3}\not=0$, the electric energy of the system can be controlled to decay sub-exponentially, at the rate of $O(t^{2}e^{-k_{3}^{2}t^{2}})$. For Bernstein modes, $k_{3}=0$, the electric energy is again controlled to be bounded, $\mathcal{E}_{\mathbf{k}}(t)\leq|\chi(\mathbf{k})|^{2}+\frac{c^{2}}{B_{0}^{2}}|\mathbf{k}_{\perp}|^{4}$, but no-decay can be said.
@@ 4 Numerical experiments
The nonlinear system is computed using a semi-Lagrangian scheme with Strang splitting [CROUSEILLES20101927, EINKEMMER2019937] for the time integration. This means from one step to the next, we first solve
@@
The code is written in JAX, and the computation was performed on an NVIDIA Quadro RTX 6000 GPU with 24 GB of memory. The data and code for results presented here are openly available in the GitHub repository [^2] https://github.com/Peiyi-wisc/Magnetized_VP_control.git, and archived in Zenodo [^3] https://doi.org/10.5281/zenodo.15109942.
@@ 4.1.1 Initial perturbation as a sine wave in space
When $c=0$, the corresponding electric energy and Fourier mode coefficient are presented in Figure 12(a)(e). The electric energy is suppressed under $10^{-5}$ for all time, and the density profile has much less variation comparing Figure 11 to Figure 10.
We also tested setting $c$ to be at different values. The results are shown in Figure 12, with $c=0.001,c=0.005,c=0.01$. For all these tests, the electric energy can be controlled below the level of $10^{-5}$ for all time, thus suppressing the instability as in Figure 9. However, we also notice that in Figure 12(d)(h), for $c$ being $0.01$, the electric energy gradually grows after $t=300$. This indicates that the perturbation induced by the term $c\frac{\sin(B_{0}t)}{B_{0}}\hat{\mu}$ in (3.10) might lead to another instability in the system.
@@ 4.2 Gaussian equilibrium
Gaussian equilibrium is expected to be a linearly stable equilibrium state according to the linear analysis when the initial perturbation is small. We now study the situation when a large perturbation is posed on the Gaussian equilibrium. This is beyond the linear stable regime, and potential instability may be observed. This is indeed the case. We set $B_{0}=1$, adopt the Kelvin-Helmholtz type instability example in [vu:hal-04016348] and use the following initial condition:
@@ 5 Conclusion
This paper studies the stabilization and control of plasma dynamics in the presence of an external magnetic field, modeled by the magnetized Vlasov–Poisson system. We derived the corresponding dispersion relation and formulated the associated Penrose-type stability condition (Proposition 2.2). As representative examples, we considered two equilibrium distributions: a Gaussian equilibrium and a ring-shaped equilibrium. The latter admits an unstable root in the dispersion relation, leading to linear instability (the Dory–Guest–Harris instability) and exponential growth of the electric energy in the linear regime.
Building on this analysis, we investigated the control problem of suppressing instability through an externally imposed electric field. The dispersion relation provides a central tool for the design of such controls. We proposed two control strategies, both based on eliminating unstable roots in the dispersion relation. One of these strategies recovers the free-streaming solution as a special case, which further enables an explicit characterization of the electric energy and a proof that it remains bounded. Numerical experiments demonstrate the effectiveness of the proposed strategies and indicate that stabilization persists beyond the strictly linear regime, including in certain nonlinear settings.
Stabilizing plasma dynamics via external fields is a key component in the broader objective of controlled fusion. While the present work is carried out in an idealized setting, it provides a step toward systematic control design in higher-dimensional kinetic models relevant to fusion applications. An important future direction is to extend the pole-elimination methodology to reduced descriptions such as drift-kinetic and gyrokinetic regimes, where the presence of multiple time and length scales introduces additional analytical and computational challenges. On the computational side, advancing to gyrokinetic scaling will likely require efficient large-scale solvers, potentially accelerated by GPU-based implementations.
@@ Acknowledgements
P.C. thanks Genia Vogman, Xiaochuan Tian, and Qihao Ye for useful discussions on the Dory-Guest-Harris instability and related details. R.J. is supported by the National Science Foundation under Grant No. 2409066. P.C., Q.L., and Y.Y. are supported by the National Science Foundation under Grant No.2308440.
@@
The authors have no competing interests to declare that are relevant to the content of this article.
@@ Appendix A Computation of $P(\mathbf{k},\lambda)$
here $M(\cdot,\cdot,\cdot)$ is the Kummer’s function (or the hypergeometric function), the last line is obtained in [integral-Gradshteyn-Ryzhik-2014] together with Kummer’s transform. One specific property of Kummer’s function is that for $j$ being nonnegative integers, $M(-j,\cdot,\cdot)$ can be expressed as a polynomial of degree no more than $j$.